### v1

# T3: A Novel Zero-shot Transfer Learning Framework Iteratively Training on an Assistant Task for a Target Task

Abstract
Long text summarization, gradually being essential for efficiently processing large volumes of information, stays challenging for Large Language Models (LLMs) such as GPT and LLaMA families because of the insufficient open-sourced training datasets and the high requirement of contextual details dealing. To address the issue, we design a novel zero-shot transfer learning framework, abbreviated as T3, to iteratively training a baseline LLM on an assistant task for the target task, where the former should own richer data resources and share structural or semantic similarity with the latter. In practice, T3 is approached to deal with the long text summarization task by utilizing question answering as the assistant task, and further validated its effectiveness on the BBC summary, NarraSum, FairytaleQA, and NLQuAD datasets, with up to nearly 14% improvement in ROUGE, 35% improvement in BLEU, and 16% improvement in Factscore compared to three baseline LLMs, demonstrating its potential for more assistant-target task combinations.
T3: A Novel Zero-shot Transfer Learning Framework Iteratively Training on an Assistant Task for a Target Task
Xindi Tong1,2, Yujin Zhu1, Shijian Fan1, Liang Xu1 1Artificial Intelligence Innovation Center, Yangtze Delta Region Institute of Tsinghua University 2Nanyang technological university to0001di@e.ntu.edu.sg, zhuyujin@tsinghua-zj.edu.cn, fanshijian@tsinghua-zj.edu.cn, Correspondence: xlpaul@126.com
1 Introduction
In recent years, long text summarization has gradually played an important role in various real-world scenarios such as current affairs commentary, book summaries, news aggregation, and literature research, benefiting from the fast developing of Large Language Models (LLMs) Nallapati et al. (2016); Vaswani et al. (2023); Dubey et al. (2024); Islam and Moushi (2024). Different from classical summarization tasks, which typically summarize texts of around 500 words or less into 100 or fewer words, long text summarization involves dealing with much longer texts. Conceivably, as the original text length increases, the amount of contextual details and relevance required in the summary also grows, bringing new challenges to LLMs.
Considering solutions, we find that long text summarization could be categorized to a more general task, which shares two crucial issues: (1) This task lacks sufficient diverse open-source datasets to post-train or fine-tune LLMs, which might be due to the task itself being relatively new and special, or because the labeling details of the task make the construction of its datasets so complex and costly, whatever through manual annotation or algorithmic generation. (2) The output of this task demands a high level of contextual details and consistence. This makes it challenging for existing prompt engineering techniques Schick and Schütze (2021); Chakraborty and Pakray (2023), most of which rely on task-oriented predefined rules to tune LLMs, thus probably losing insight into richer contextual information, as well as failing to capture similar semantic laws behind different tasks.
In this paper, we propose a novel zero-shot Transfer learning framework iteratively training on an assistant Task for a target Task (abbreviated as T3). T3 fine-tunes the chosen baseline LLM through the assistant task that must meet two basic criteria: (1) It should have relatively abundant open-source datasets to address the first "insufficient dataset" issue. (2) It should exhibit similarity at the structural or semantic level to the target task to address the second "more details" issue mentioned above. Moreover, inspired by the ability of emergence from the new generation of LLMs to understand more complex semantics, an iterative optimization strategy is adopted to enable the used LLM to learn deeper from the assistant task and summarize useful rules (called as "experiences" in this paper) automatically by itself. Hereafter, the fine-tuned LLM is applied to the target task in a zero-shot manner. Further in practice, with long text summarization as the target task, we find that the Question Answering task (QA) is suitable as the assistant task because it has various open-source datasets, its used text is relatively long, and its question-answer pairs naturally contain richer entities and relationships. In addition, the Question Generation task (QG) is used during the training process of T3 to help LLMs better understand common and different contextual features between the assistant and the target tasks. Our contributions are concluded as follows:
  • $\bullet$
    We propose a novel zero-shot transfer learning framework named T3. Different from most previous works, the baseline model in T3 only learns from the assistant task and then directly runs for the target task in zero-shot way. Further, T3 facilitates the transfer of knowledge from QA and QG tasks to the long text summarization task.
  • $\bullet$
    We adopt an iteration strategy in training process to make the used baseline model conclude useful rules automatically, with specially designed prompts and selected metrics, in order to ensure the high quality of the summaries generated during the test process in T3.
  • $\bullet$
    We demonstrate significant effectiveness of T3 on four datasets across different tasks by comparing with seven representative baseline LLMs.
2 Related Work
2.1 Summarization
Early automatic text summarization has begun with Luhn’s keyword frequency method and Edmundson’s positional weighting in the 1950s. Machine learning approaches have emerged in the 1990s, such as Kupiec’s Bayesian classifier. Later, deep learning methods like TextRank Mihalcea and Tarau (2004) advance the field. Recent research focuses on improving faithfulness in abstractive summarization, with PEGASUS Zhang et al. (2020) and methods like ECC Zhang et al. (2022b). Other works include contrastive learning Chen et al. (2021), graph attention Zhu et al. (2021), and few-shot approaches Schick and Schütze (2021). LLMs, like GPT, also contribute to faithful summarization Zhang et al. (2023).
Texts with more than 1,000 words could be considered as long text Tuteja and González Juclà (2023). Transformer-based models are efficient to handling of lengthy texts. Longformer Beltagy et al. (2020) uses a local attention mechanism to reduce computational complexity, enabling efficient processing of longer texts without losing global context. HyperGraph Transformers introduced by Zhang et al. (2022a) tackle the complexity and context retention issues in Transformer-based models. Additionally, datasets like SQuALITY are crucial for advancing long-document summarization Wang et al. (2022). GPT could also applied for long text summarization which shows better performance than LLaMA Fan et al. (2024).
2.2 Question Answering
Datasets like NQ Kwiatkowski et al. (2019) and SQuAD Rajpurkar et al. (2016, 2018) have been foundational for the task. NQ focuses on real-world queries with Wikipedia annotations, while SQuAD emphasizes extractive QA. MS MARCO Bajaj et al. (2016) and DuReader He et al. (2018) have broaden the QA field with free-form answers and non-English tasks. Recent works like Eo et al. (2023) enhance QA diversity via multimodal inputs and iterative generation. Transformer-based models like BERT Devlin et al. (2019), fine-tuned on NQ and SQuAD, improve QA precision and recall. Further, Eo et al. (2023) distills complex information for summarization, aiding coherent QA generation. Here, QA is used as an assistant task for the proposed T3 on the summarization task.
2.3 Question Generation
Earlier work by Du et al. (2017) has highlighten the effectiveness of sequence-to-sequence models for QG in reading comprehension, while Song et al. (2018) improve QG accuracy using a multi-perspective context-matching model. Chali and Baghaee (2018) explore summarization for opinion-based QG to reduce repetition. Lyu et al. (2021) introduce an unsupervised approach, utilizing summaries to generate diverse questions with less lexical overlap. Frameworks like MQAG by Manakul et al. (2023) ensure coherence in QG through summarization. Recently, Choudhary and Du (2024) has proposed the QAEVENT paradigm, using summarization to represent document-level events as question-answer pairs. In this paper, QG is considered as a generation strategy in the training process of T3.
Figure 1: General workflow of T3
2.4 Zero-shot Transfer Learning
Recent advancements in zero-shot transfer learning have significantly improved the capabilities of LLMs across diverse tasks to solve problems such as data misalignment, label mismatches, and the limitations of pre-trained models without further fine-tuning Kojima et al. (2023); Zhao et al. (2023). Furthermore, multilingual LLMs have demonstrated impressive zero-shot learning abilities in multimodal settings and cross-lingual instruction tuning, showing that LLMs trained on one language can generalize to others Hu et al. (2023). ARL2 introduces an adaptive retriever learning technique that aligns with LLMs for better zero-shot generalization, particularly in tasks like QA Zhang et al. (2024). ZeroG applying GPT models focuses on zero-shot transfer for graph-based tasks, pioneering cross-dataset node classification by combining graph representations with LLM Li et al. (2024). In this paper, a novel zero-shot transfer learning strategy using both assistant and target task is proposed.
3 Methodology
In this section, we first demonstrate the general workflow of proposed T3, and then formulate the summarization task under T3, thereafter dive into the detailed training and test processes.
3.1 T3 General Workflow
The general workflow of T3 is shown in Figure 1. The keypoint of T3 is that: in the whole training process, both the assistant task and the target task share the samples of the assistant task. At first, the used LLM  [^1] GPT-4o is recommended, while any other large models with sufficient parameter capacity to possess emergence could be considered. acquires experience to deal with the assistant task through the labeled data, which means a deeper comprehension of the samples. Subsequently, LLM iteratively tries to accomplish the target task on these same samples and gain the experience for the target task by leveraging the previous learned contextual experiences, as well as being modulated by the predefined conditions. Finally, the model applies the experiences required for both the assistant and target tasks during the test process.
3.2 T3 for Summarization
Now the general workflow of T3 is implemented to our target task: summarization, with the question-answering issue as the assistant task.
Figure 2: Training and test process of T3 for summarization task.
3.2.1 Task Formulation
Given the training dataset for Question-Answering task (QA), each sample from the dataset contains both the text $D_{i}$ and several labeled question-answer pairs $QA_{i}$ (where $i$ implies the $i$th sample), which form the text set $D=\{D_{1},D_{2},\ldots,D_{n}\}$ and the label set $QA=\{QA_{1},QA_{2},\ldots,QA_{n}\}$. Relatively, the question-answer pairs generated by LLM from $D_{i}$ are written as $QA_{i}^{\prime}$. Given the test dataset for Summarization task (Sum), each sample from the dataset contains a text $D_{j}$, while LLM needs to generate the summary of the text $Sum_{j}^{\prime}$. In training process, the prompts $P_{QA}$ and $P_{Sum}$ (see Appendix A.2 and A.3) are used for LLM to generate the question-answer pairs $QA_{i}^{\prime}$ and the summary $Sum_{i}^{\prime}$, respectively. At the same time, LLM learns experiences from both tasks and records them in the form of several rules into $Exp_{QA}$ and $Exp_{Sum}$.
Figure 3: Test process of T3 for summarization task on QA dataset.
3.2.2 Training Process
The overview of T3 for summarization is illustrated in Figure 2. The training process can be found at the top of the figure, while the test process is demonstrated at the bottom. Both inputs and outputs in these processes are written in bold in the figure. Furthermore, the detailed information of the training process is given in Algorithm 1 in Appendix A.1, including the following key steps:
Updating $Exp_{QA}$: At first, both summary and QA generation experiences $Exp_{Sum}$ and $Exp_{QA}$ are initialized as $null$. Then, LLM generates new QA pairs $QA_{i}^{\prime}$ for each text $D_{i}$ and updates $Exp_{QA}$ according to several rules concluded by itself through comparing $QA_{i}\textquoteright$ ’ with the ground truth $QA_{i}$. The prompt $P_{QA}$ is predefined to control all steps mentioned in this part.
Updating $Exp_{Sum}$: In this part, LLM follows the predefined prompts $P_{Sum}$ and try to summarize the text $D_{i}$. As an prior-knowledge that might bring extra contextual information, both $QA_{i}$ and $QA_{i}\textquoteright$ ’ are also taken into account as the inputs. After the current summary $Sum_{i}^{\prime}$ is generated, three thresholds are used to investigate whether the generated content is well enough. LLM would repeat the summary generation step and update its experience $Exp_{Sum}$ every iteration, till the stop conditions are fulfilled.
Thresholds Used in Iteration: The iterative process mentioned above continues until the generated summary $Sum_{i}^{\prime}$ simultaneously meets the required thresholds for similarity, readability Flesch (1948), and compression rate (or the number of iterations reach the preset maximum iteration times $K$). In details, readability R is calculated as:
$\text{R}=206.835-1.015\times\left(\frac{TW}{TSE}\right)-84.6\times\left(\frac{TSY}{TW}\right)$ (1)
where, $TW$ is the total number of words in the given text, $TSE$ is the total number of sentences, and $TSY$ is the total number of syllables. $TW/TSE$ calculates the average sentence length: a higher average sentence length makes the text harder to read, so it subtracts from the score. $TSY/TW$ calculates the average syllables per word: more syllables per word reduce the score because longer words tend to make the text more difficult to read. Moreover, the compression rate C is calculated as:
$\text{C}=\text{length}(Summary)/\text{length}(Text)$ (2)
which is used to control the length of the generated summary.
3.2.3 Test Process on Summarization Dataset
According to the bottom of Figure 2, during the test process, LLM would first use $Exp_{QA}$ learned from the training process to generated auxiliary question-answer pairs, and bring them with the updated $Exp_{Sum}$ to generate the final summary. Prompts in test process are given in Appendix A.4.
3.2.4 Test Process on QA Dataset
In particular, as known from Figure 2. We design QG, a more challenging text generation strategy for LLM, to replace the text understanding strategy of predicting answers for the questions from the given QA dataset during the training process, in order to better fit LLM to the summarization task. A by-product of this approach is that we can even run the QA dataset in test process to generate summaries for its own texts based on its QA pairs, and adopt appropriate metrics for effectiveness evaluation, as shown in Figure 3. Prompts in test process for QA dataset are given in Appendix A.5.
4 Experiments
This section presents experimental details on four datasets, including two narrative dataset and two news datasets. We compare the quality of the summaries generated by the baseline models and our model using the generation algorithm from different aspects and perform an ablation study to analyze the influence of different components of T3.
4.1 Settings
4.1.1 Datasets
Four datasets are used, including two summarization ones (BBC summary and NarraSum), and two QA ones (NLQuAD and FairytaleQA):
BBC Summary includes 2,225 BBC news from 2004-2005, used for categorization and summary research Greene and Cunningham (2006).
NarraSum has 122K narrative synopses from movies and TV episodes, aimed at improving summarization research Zhao et al. (2022).
NLQuAD contains 31,000 questions and answers derived from 13,000 BBC news, for non-factoid long QA, focusing on multi-sentence descriptive answers to enhance document-level language models and challenges models with long context and complex language understanding Soleimani et al. (2021).
FairytaleQA includes 278 children’s stories from Project Gutenberg, with questions and answers created by educational experts, aimed at assessing and training narrative comprehension for students (K-8) and models Xu et al. (2022).
In experiments, 386 long articles from BBC summary, 260 long stories from FairytaleQA, 100 long stories from NarraSum, 501 long articles from NLQuAD are randomly picked up for evaluation. Each selected texts has a median word count exceeding 1,000.
4.1.2 Evaluation Metrics
To make more comprehensive evaluation, especially focusing on generated text quality and hallucinations alleviation, a range of classical and innovative metrics are used here:
ROUGE compares the overlap of n-grams, word sequences, and word pairs between the generated summary and reference summaries Lin (2004).
BLEU compares the overlap of n-grams between the generated text and reference translations, with higher scores indicating closer alignment with human translations and thus better quality Papineni et al. (2002).
FactSumm FactSum uses named-entity recognition and relation extraction to extract facts from the source text and the generated summary to evaluate fact consistency by comparing extracted facts represented as triples (subject, relation, object) Zhang et al. (2021).
ChatGPT evaluates factual inconsistency in summaries, outperforming previous methods with fewer hallucinations Luo et al. (2023).
4.1.3 Baseline Models
Seven advanced LLMs from three most representative series are selected as the baselines. All of them are accessed through Application Programming Interface (API), which not only standardizes the experiment for reproducibility and consistency, but also saves computational resources.
GPT Series are developed by OpenAI and show strong language comprehension and generation capabilities for various natural language processing tasks Achiam et al. (2023). Here, GPT-3.5-turbo, GPT-4 Turbo, and the newest lauched GPT-4o are adopted here.
Claude Series introduce advanced multimodal capabilities developed by Anthropic Anthropic (2023b). Claude 3.5 Sonnet is the newest one of them, showing strong performance in graduate-level and text-based reasoning Anthropic (2023a). Claude 3.5 Sonnet and Claude 3.0 Sonnet are chosen here.
Gemini Series surpasses Google’s capabilities in understanding images, audio, video, and text Team et al. (2023). Currently, Gemini-1.5-pro is the newest and best-performing one of them Reid et al. (2024). Both Gemini-1.5-pro and Gemini-1.0-pro are used here.
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
BBC summary
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.4019 | 0.4171 0.1814 | 0.2145 0.3697 | 0.3944 0.1086 | 0.1094 0.6364 | 0.7012
GPT-4-turbo 0.3733 | 0.3683 0.1476 | 0.1414 0.3373 | 0.3327 0.0851 | 0.0769 0.6091 | 0.6901
GPT-3.5-turbo 0.3936 | 0.3788 0.1775 | 0.1695 0.3604 | 0.3478 0.1050 | 0.0858 0.5752 | 0.4780
Gemini-1.5-pro 0.3362 | 0.3510 0.1148 | 0.1258 0.3036 | 0.3169 0.0505 | 0.0647 0.2588 | 0.6049
Gemini-1.0-pro 0.3490 | 0.3358 0.1280 | 0.1392 0.3175 | 0.3117 0.0599 | 0.0722 0.3576 | 0.5561
Claude-3.5-sonnet 0.3908 | 0.4354 0.1659 | 0.2113 0.3588 | 0.4036 0.0970 | 0.1300 0.6343 | 0.6583
Claude-3.0-sonnet 0.4123 | 0.4090 0.1826 | 0.1808 0.3778 | 0.3739 0.1108 | 0.1074 0.6002 | 0.5043
NarraSum
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.3114 | 0.2886 0.0731 | 0.0858 0.2844 | 0.2566 0.0268 | 0.0255 0.9775 | 0.9969
GPT-4-turbo 0.2914 | 0.2727 0.0772 | 0.0742 0.2544 | 0.2445 0.0336 | 0.0360 0.9775 | 0.9744
GPT-3.5-turbo 0.2860 | 0.2528 0.0752 | 0.0588 0.2650 | 0.2300 0.0279 | 0.0215 0.9520 | 0.9591
Gemini-1.5-pro 0.2599| 0.2695 0.0521 | 0.0539 0.2500 | 0.2532 0.0175 | 0.0199 0.9326 | 0.9357
Gemini-1.0-pro 0.2897 | 0.2825 0.0626 | 0.0574 0.2545 | 0.2636 0.0165 | 0.0209 0.9285 | 0.9122
Claude-3.5-sonnet 0.2919 | 0.2978 0.0717 | 0.0779 0.2673 | 0.2723 0.0310 | 0.0323 0.9704 | 0.9814
Claude-3.0-sonnet 0.2990 | 0.2587 0.0675 | 0.0547 0.2645 | 0.2251 0.0274 | 0.0205 0.9469 | 0.9561
NLQuAD
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.2437 | 0.3100 0.0848 | 0.1303 0.2271 | 0.2927 0.0033 | 0.0142 0.5981 | 0.6395
GPT-4-turbo 0.2419 | 0.2366 0.0739 | 0.0736 0.2216 | 0.2162 0.0044 | 0.0036 0.5333 | 0.4916
GPT-3.5-turbo 0.2150 | 0.2221 0.0775 | 0.0830 0.2007 | 0.2069 0.0027 | 0.0027 0.5530 | 0.4464
Gemini-1.5-pro 0.2299 | 0.2419 0.0697 | 0.0797 0.2134 | 0.2244 0.0025 | 0.0047 0.5085 | 0.6509
Gemini-1.0-pro 0.2626 | 0.2811 0.0929 | 0.1105 0.2455 | 0.2648 0.0054 | 0.0125 0.6457 | 0.7164
Claude-3.5-sonnet 0.3203 | 0.4017 0.1280 | 0.1874 0.3006 | 0.3792 0.0235 | 0.0500 0.6571 | 0.8276
Claude-3.0-sonnet 0.3235 | 0.3372 0.1270 | 0.1327 0.3016 | 0.3122 0.0197 | 0.0237 0.5889 | 0.6169
FairytaleQA
w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
GPT-4o 0.1334 | 0.1486 0.0315 | 0.0372 0.1234 | 0.1378 0.0004 | 0.0007 0.9855 | 0.9892
GPT-4-turbo 0.1468 | 0.1414 0.0351 | 0.0336 0.1345 | 0.1284 0.0009 | 0.0010 0.9744 | 0.9767
GPT-3.5-turbo 0.1222 | 0.1223 0.0298 | 0.0307 0.1131 | 0.1074 0.0003 | 0.0004 0.9982 | 0.9737
Gemini-1.5-pro 0.1540 | 0.1756 0.0380 | 0.0434 0.1430 | 0.1546 0.0010 | 0.0013 0.9723 | 0.9853
Gemini-1.0-pro 0.1414 | 0.1343 0.0325 | 0.0243 0.1422 | 0.1339 0.0003 | 0.0014 0.9784 | 0.9424
Claude-3.5-sonnet 0.1597 | 0.2096 0.0373 | 0.0607 0.1494 | 0.1952 0.0017 | 0.0047 0.9956 | 0.9997
Claude-3.0-sonnet 0.1720 | 0.1891 0.0408 | 0.0490 0.1581 | 0.1735 0.0017 | 0.0033 0.9747 | 0.9892
Table 1: Experimental Results on BBC summary, NarraSum, NLQuAD and FairytaleQA datasets where each LLM without (left col) or with (right col) T3.
4.2 Implementation Details
The comparative experiment are designed and executed to validate both the feasibility and effectiveness of the proposed T3, by investigating the generated summarization results for all four datasets from all used LLMs with/without using T3.
The baseline summary of each text for each LLM is generated by using original text and the following prompt: "You are one helpful text assistant. Based on the input text, provide a summary in proper length". Correspondently, each LLM with T3 is guided by the experience and prompts derived from training process.
For BBC summary and NarraSum in test process, both the original texts from them, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
For FairytaleQA and NLQuAD in test process, 10 texts and their corresponding QA pairs from each dataset are selected for training, while the remaining texts are used for testing. Both the original texts from the datasets, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
The higher scores from the used metrics are, indicating the better generation of summaries. It should be noted that the metric Factscore was calculated by GPT-4o for NarraSum and FairytaleQA, while calculated by FactSumm for BBC summary and NLQuAD, because it cannot extract facts from the story source text in practice so that does not perform well on narratives.
For BBC summary and NarraSum, the provided summary was used as the reference, while for FairytaleQA and NLQuAD datasets, the original text was used as the reference. As mentioned in Part 3.2.4, T3 can also deal with QA datasets for summarization task. When the dataset does not provide reference summaries, the generated summary can be directly compared to the original text through certain automated metrics tools (e.g., ROUGE) Louis and Nenkova (2009). According to Saggion and Poibeau (2013), the overlap of the generated summaries with the original text is important in assessing information coverage and the validity.
Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Readbility Factscore
Claude-3.5-sonnet BBC summary
Full T3 0.4354 0.2113 0.4036 0.1300 54.71 0.6583
w/o summary experience 0.4026 0.1683 0.3599 0.1331 55.08 0.5832
w/o qa experience 0.4155 0.2072 0.4005 0.3190 55.64 0.5857
GPT-4o NLQuAD
Full T3 0.3100 0.1303 0.2927 0.0142 44.53 0.6395
w/o summary experience 0.3628 0.1529 0.3402 0.0309 42.57 0.7785
w/o qa experience 0.4116 0.2029 0.3899 0.0531 45.77 0.6523
Table 2: Ablation study results on BBC summary and NLQuAD datasets.
For models from GPT or Claude series, default values for temperature and other related hyper-parameters are applied. For Gemini models, we set ’HARASSMENT’, ’HATE_SPEECH’, ’SEXUALLY_EXPLICIT’, ’DANGEROUS’ in safety setting into ’block_none’.
4.3 Results
Table 1 illustrates the performance of all used models with (right "T3" column) and without (left "w/o" column) T3 on all adopted datasets. The better one of the two values from each row of the two columns is marked in bold. Especially, for each row where all results from columns with T3 are better than those columns without T3, the row is colored blue. From this table, we observe that:
(1) LLMs using T3 outperforms themselves without T3 under most metrics of most datasets, which demonstates the feasibility and effectiveness of T3. Specially, the average improvement rates of GPT-4o, Gemini-1.5-pro, and Claude-3.5-sonnet using T3 compared to the baselines without T3 across different tasks are calculated and shown that these LLMs with T3 nearly improves the ROUGE score by 14%, the BLEU score by 35%, and the Factscore score by 16%.
(2) Of all the models in the four series, it is almost always the newest ones with T3 that achieve the most significant results comparing with itself without T3. For example, GPT-4o is 4% and 5% higher on Rouge-1 metrics compared to GPT-4 and GPT-3.5 on BBC Summary, respectively. These results could be due to the improvement of capabilities for latest LLMs on handling longer texts and more comprehensive tasks through understanding more complicated prompts and locating better details and events within lengthy texts Anthropic (2023a); Islam and Moushi (2024); Reid et al. (2024).
(3) Four used datasets are divided into news- and narrative-styles. For most models with/without T3, the generated summaries on narrative-style datasets NarraSum seems not as good as that of news-style ones BBC Summary. According to Genette (1980); Van Dijk (2013); Barthes (2016), the structure of news-style texts is compared with that of narrative-style texts, highlighting how news is easier to be summarized because of its concise layout and fixed pattern, while the narrative texts is more complexity, including chronological order, shifts in point of view, and narrative hierarchy, so that increasing the difficulty of summarization.
(4) The improvement of LLMs with T3 seems more significant on NLQuAD and FairytaleQA, i.e., QA task-oriented datasets, which might be benefited from using the prior information of questions and answers pairs to enhance the contextual analyzing capability of T3. Moreover, incorporating QA information implies to improve summarization quality by enhancing text comprehension, focusing on key information, increasing coverage, ensuring coherence, and aligning summaries with user needs Fabbri et al. (2022); Siledar et al. (2024).
4.4 Ablation Study
We further analyze the influence of different components of T3. Table 2 shows the experimental results for Claude-3.5-sonnet after removing summary experience and QA generation experience on the BBC summary dataset, and the results for GPT-4o after removing summary experience and QA generation experience on the NLQuAD dataset. For models with less reasoning ability, longer inputs and complex instructions of T3 might diminish their performance, while for more advanced models, each component of T3 could be utilized to enhance their overall effectiveness.
5 Conclusion
A novel T3 framework is proposed to address the challenge of long text summarization, using QA datasets to iteratively train baseline model at first, then to deal with summarization in a zero-shot way. Experiments on four datasets among seven representative baselines under five evaluation metrics validate the significant effectiveness of T3. Since T3 is model-agnostic and theoritically could be used for any suitable assistant task and target task pair, the future work will aim to explore the potential of it in a broader range of task combinations.
Limitations
When using the proposed T3 framework in real scenarios, the following limitations that need to be further optimized:
Limited Baseline Models Due to the costs and time constraints of API, we only test T3 on models from GPT, Claude, and Gemini series. More open-source models (such as LLaMA) and their private deployment situation will be considered in the future.
Limited Tasks Few empirical guidelines for the selection of assistant tasks are given in this paper. In the future, the matching rules of assistant tasks and target tasks will be comprehensively explored in various task combinations.
Limited Evaluation Metrics As to evaluate generation results, BLEU and ROUGE focus on word overlap and neglect fluency, semantics, and factual accuracy, while FactSumm aims to consistency but struggles with abstract concepts. For summarization task, more suitable evaluation metrics should be discussed.
Ethics Statement
Data Availability and Safety The training and test data used for summarization and QA in this paper are publicly available. While the original datasets were filtered, some content may still include sensitive material, such as reports on violent crimes or incidents. Furthermore, when utilizing the Gemini model service, the safety mode is set to ’block_none,’ which may result in outputs containing violent or explicit content.
Usage of Large Language Models Models like GPT, Gemini, and Claude are utilized to generate QA results, refine prompt descriptions, and produce text summaries for input texts in summarization tasks. The generated text is used exclusively for experimental and analytical purposes, and the findings are presented in the relevant sections.
Appendix A Appendix
A.1 Training Steps of T3 for Summarization
Algorithm 1 T3 Training for Summarization
Input: Text set $D=\{D_{1},D_{2},\ldots,D_{n}\}$ and label set $QA=\{QA_{1},QA_{2},\ldots,QA_{n}\}$ of the assistant QA task; Initialized experience $Exp_{QA}$ and $Exp_{Sum}$ for the QA and Summarization tasks, respectively.
Parameter: Predefined prompt $P_{QA}$ and $P_{Sum}$ for QA and Summary generation, respectively. Maximum iteration for each text $K$. Similarity $S$, readability $R$, and compression rate $C$ thresholds between the generated summary and the original text, respectively. The used baseline LLM $M$.
Output: Updated QA generation experience $Exp_{QA}$. Updated summary generation experience $Exp_{Sum}$
1:  Initialize $Exp_{QA}\leftarrow\emptyset$, $Exp_{Sum}\leftarrow\emptyset$
2:  for $i=1$ to $n$ do
3:     Generate QA pairs $QA_{i}^{\prime}$ for $D_{i}$ using $Exp_{QA}$ based on $P_{QA}$ and model $M$.
4:     Learn from the generation process and update $Exp_{QA}$ using model $M$.
5:     for $k=1$ to $K$ do
6:        Generate summary $Sum_{i}^{\prime}$ for $D_{i}$ using $Exp_{Sum}$ based on $P_{Sum}$, $QA_{i}^{\prime}$, $QA_{i}$, and model $M$.
7:        Learn from the generation process and update $Exp_{Sum}$ using model $M$.
8:        Calculate the similarity between $Sum_{i}^{\prime}$ and $D_{i}$ using cosine similarity $S_{i}$.
9:        Calculate the readability $R_{i}$ of $Sum_{i}^{\prime}$ based on equation 1.
10:        Calculate the compression rate $C_{i}$ of the generated summary based on equation 2.
11:        if $S_{i}>S$ and $R_{i}>R$ and $C_{i}<C$ then
12:           break
13:        end if
14:     end for
15:  end for
A.2 Prompt for QA Generation and Experience Updating in Training Process
You are one helpful text assistant and you are good at learning new experience. First please help us to do the following:
1. Based on input text, provided questions and QA pair generation experience, generate questions and corresponding answers to original text that can reflect all the subjects and all the key points of input article. The format of the generated question pairs should be same as the provided question pairs.
2. Update the learning experience of how to generate better question pairs during the question generation process by comparing the generated questions and original question pairs based on the original text or other methods. You need to consider all the key points to make a good QA pairs and the learning experience should be general. DO NOT INCLUDE DETAIL EXAMPLES BUT PROVIDE rules AS DETAILED AS POSSIBLE.DO NOT EXCEED 600 WORDS.
The format of the output is as follows:
{ "Generated_QA_pairs":
{ "1": { "Question": "xxxx", "Answer": "xxxx"},
"2": { "Question": "xxxx", "Answer": "xxxx"},
"QA_generation_experience": "Answer point by point text in one line"}
A.3 Prompt for Summarization Generation and Experience Updating in Training Process
You are one helpful text assistant. Please help us to do the following:
1. Based on input text, provided questions, generated QA pairs and summary generation experience, generate summary corresponding answers to original text that can reflect all the subjects and all the key points of input article in proper length.
2. Update the summary generation experience of how to generate better summary with the help of provided questions first by the generated summary and all provided questions (the generated and the original) and then by what you learn when generating questions in question generation experience.
Here is the target for the generated summary:
THE GENERATED SUMMARY SHOULD HAS HIGHER ROUGE AND BLEU SCORES ALIGNED WITH THE ORIGINAL TEXT THAN THE DIRECTED GENERATED ONES. The summary should include essential terms, phrases, and sentences that closely mirror the original text to ensure high precision and recall. Notice Aim to maximize the overlap with the reference summary by closely following the wording, sentence structure, and style of the original text. Ensure that the summary is concise but comprehensive, avoiding unnecessary details or deviations. You need also consider the fluency of summary.
You need to consider all the key points to make a good summary and the learning experience should be general.(for example, like does not lose the detail in the story and pay attention to the connection of sentences, listing some point to improve the accuracy and the fluency of the story, YOU STILL NEED TO LIST OTHER rules THAT MAY IMPORTANT,Every point should as detailed as possible).
You can also think some methods to ensure a better summary, such as voting.
Answer in one line and point by point. DO NOT EXCEED 600 WORDS.’ ’The format of the output is as follows:
{ "Summary": "text in one line",
"Summary_generation_experience": "Answer point by point text in one line"}.
You must follow this format.
A.4 Prompt in Test Process on Summarization Dataset
You are a helpful text assistant. Please help us to do the following:
Based on the input text, with the help of QA generation rules and summary rules, here are the steps to do the summary:
First, generate QA pairs based on the generation rules.
Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
Article: [Article]
Summary Generation Experience: [Summary generation experience]
QA Generation Experience: [QA generation experience]
A.5 Prompt in Test Process on QA Dataset
You are a helpful text assistant. Please help us to do the following:
Based on the input text and the questions given, with the help of QA generation rules and summary rules, here are the steps to do the summary:
First, generate QA pairs based on the generation rules.
Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
Article: [Article]
Question Pair: [Question Pair with answer]
Summary Generation Experience: [Summary generation experience]
QA Generation Experience: [QA generation experience]
A.6 Summary Experience for News-style Datasets
1. Ensure all key points such as the formation, evolution, and current status of Boko Haram, and the contrasting views on security are included.
2. Maintain original wording and sentence structure for precision and recall.
3. Avoid unnecessary details, focusing on core information for conciseness.
4. Ensure logical flow and fluency in the summary.
5. Use specific terms and phrases from the text to maximize overlap with the reference summary.
6. Consider using a voting system to select the best summary from multiple drafts.
7. Verify the summary reflects the original text accurately without speculative information.
8. Address distinct key points or subjects in each question for comprehensive coverage.
9. Compare generated questions with original ones to ensure all critical aspects are covered.
10. Ensure answers are directly supported by the text, avoiding speculative or inferred information.
A.7 Summary Experience for Narratives-style Datasets
1. Ensure all key points and characters from the text are covered to avoid missing crucial details,use Details to Enhance Emotional Impact such as Incorporate symbolic details.
2. Maintain the original sequence of events to reflect the narrative flow accurately.
3. Use language and phrases closely mirroring the original text for higher precision and recall.
4. Keep the summary concise but comprehensive, avoiding unnecessary details while ensuring all significant events are included.
5. Focus on fluency, ensuring the summary reads smoothly and logically. Preserve the logical flow and connection between events for coherence.
6. Review and refine the summary multiple times to maximize overlap with the reference summary.
7. Pay attention to the connection of sentences to maintain coherence.
8. Include essential terms and phrases to ensure high precision and recall.
9. Emphasize Themes and Symbolism: Don’t just recount the plot; also highlight the underlying themes or symbolic meanings.
10. Aim to maximize ROUGE and BLEU scores by closely following the wording, sentence structure, and style of the original text.
11. Highlight the emotional tones and themes present in the original text. This adds depth to the summary and helps convey the underlying messages or morals of the story. It’s important not just to recount events but also to capture the feelings and ideas that these events evoke.
12. Ensure that no important details are lost that contribute to the storyś moral or lesson. These details are often critical to understanding the overall message of the text, so they should be preserved to maintain the integrity of the original story.
13. Consider using a voting method to select the best summary from multiple versions, such as generate 3 versions and find the best one.
A.8 QA Generation Experience for News-style Datasets
1. Ensure each question targets a distinct key point or subject from the original text.
2. Questions should be specific to elicit detailed answers that reflect core information.
3. Avoid broad or vague questions to prevent ambiguous answers.
4. Compare generated questions with original ones to ensure all critical aspects are covered, and questions are clear and concise.
5. Answers should be directly supported by the text and avoid speculative or inferred information. This approach ensures comprehensive and accurate QA pairs.
A.9 QA Generation Experience for Narratives-style Datasets
1.Ensure each question addresses a unique key point or detail to avoid redundancy. Cover all significant events, characters, and details for comprehensive coverage.
2. Keep questions clear and specific to avoid ambiguity and ensure they can be answered directly from the text.
3. Include a mix of factual, inferential, and interpretative questions to engage different levels of understanding.
4. Review the text multiple times to ensure no key points are missed and questions flow logically.
5. Pay attention to character motivations, actions, and consequences to generate deeper questions.
6. Consider the context and sequence of events to create questions reflecting the narrativeś progression.
7. Ensure answers are concise and directly derived from the text for accuracy.
8. Compare generated questions with existing ones to identify gaps or overlaps and refine accordingly.
9. Aim to create questions that encourage critical thinking and deeper engagement with the text.’

### v3

@@
 1 Introduction
-In recent years, long text summarization has gradually played an important role in various real-world scenarios such as current affairs commentary, book summaries, news aggregation, and literature research, benefiting from the fast developing of Large Language Models (LLMs) Nallapati et al.
+In recent years, long text summarization has gradually played an important role in various real-world scenarios such as current affairs commentary, book summaries, news aggregation, and literature research, benefiting from the fast development of Large Language Models (LLMs) Nallapati et al.
 (2016); Vaswani et al.
@@
 (2024); Islam and Moushi (2024).
-Different from classical summarization tasks, which typically summarize texts of around 500 words or less into 100 or fewer words, long text summarization involves dealing with much longer texts.
-Conceivably, as the original text length increases, the amount of contextual details and relevance required in the summary also grows, bringing new challenges to LLMs.
+Unlike classical summarization tasks, which typically summarize texts of around 500 words or less into 100 or fewer words, long text summarization involves dealing with much longer texts.
+Clearly, as the original text length increases, the amount of contextual details and relevance required in the summary also grows, bringing new challenges to LLMs.
 Considering solutions, we find that long text summarization could be categorized to a more general task, which shares two crucial issues: (1) This task lacks sufficient diverse open-source datasets to post-train or fine-tune LLMs, which might be due to the task itself being relatively new and special, or because the labeling details of the task make the construction of its datasets so complex and costly, whatever through manual annotation or algorithmic generation.
@@
 In this paper, QG is considered as a generation strategy in the training process of T3.
-Figure 1: General workflow of T3
 2.4 Zero-shot Transfer Learning
@@
 In this paper, a novel zero-shot transfer learning strategy using both assistant and target task is proposed.
+Figure 1: Training and test process of T3 for summarization task.
 3 Methodology
-In this section, we first demonstrate the general workflow of proposed T3, and then formulate the summarization task under T3, thereafter dive into the detailed training and test processes.
-3.1 T3 General Workflow
-The general workflow of T3 is shown in Figure 1.
-The keypoint of T3 is that: in the whole training process, both the assistant task and the target task share the samples of the assistant task.
-At first, the used LLM  [^1] GPT-4o is recommended, while any other large models with sufficient parameter capacity to possess emergence could be considered.
-acquires experience to deal with the assistant task through the labeled data, which means a deeper comprehension of the samples.
-Subsequently, LLM iteratively tries to accomplish the target task on these same samples and gain the experience for the target task by leveraging the previous learned contextual experiences, as well as being modulated by the predefined conditions.
-Finally, the model applies the experiences required for both the assistant and target tasks during the test process.
-3.2 T3 for Summarization
-Now the general workflow of T3 is implemented to our target task: summarization, with the question-answering issue as the assistant task.
-Figure 2: Training and test process of T3 for summarization task.
-3.2.1 Task Formulation
+In this section, we first formulate the summarization task under T3, and dive into the detailed training and test processes. Then, we conclude the general form of the T3 workflow.
+3.1 Task Formulation
 Given the training dataset for Question-Answering task (QA), each sample from the dataset contains both the text $D_{i}$ and several labeled question-answer pairs $QA_{i}$ (where $i$ implies the $i$th sample), which form the text set $D=\{D_{1},D_{2},\ldots,D_{n}\}$ and the label set $QA=\{QA_{1},QA_{2},\ldots,QA_{n}\}$.
@@
 At the same time, LLM learns experiences from both tasks and records them in the form of several rules into $Exp_{QA}$ and $Exp_{Sum}$.
-Figure 3: Test process of T3 for summarization task on QA dataset.
-3.2.2 Training Process
+3.2 Training Process
 The overview of T3 for summarization is illustrated in Figure 1.
@@
 Both inputs and outputs in these processes are written in bold in the figure.
-Furthermore, the detailed information of the training process is given in Algorithm 1 in Appendix A.1, including the following key steps:
-Updating $Exp_{QA}$: At first, both summary and QA generation experiences $Exp_{Sum}$ and $Exp_{QA}$ are initialized as $null$.
-Then, LLM generates new QA pairs $QA_{i}^{\prime}$ for each text $D_{i}$ and updates $Exp_{QA}$ according to several rules concluded by itself through comparing $QA_{i}\textquoteright$ ’ with the ground truth $QA_{i}$.
-The prompt $P_{QA}$ is predefined to control all steps mentioned in this part.
-Updating $Exp_{Sum}$: In this part, LLM follows the predefined prompts $P_{Sum}$ and try to summarize the text $D_{i}$.
-As an prior-knowledge that might bring extra contextual information, both $QA_{i}$ and $QA_{i}\textquoteright$ ’ are also taken into account as the inputs.
-After the current summary $Sum_{i}^{\prime}$ is generated, three thresholds are used to investigate whether the generated content is well enough.
-LLM would repeat the summary generation step and update its experience $Exp_{Sum}$ every iteration, till the stop conditions are fulfilled.
-Thresholds Used in Iteration: The iterative process mentioned above continues until the generated summary $Sum_{i}^{\prime}$ simultaneously meets the required thresholds for similarity, readability Flesch (1948), and compression rate (or the number of iterations reach the preset maximum iteration times $K$).
-In details, readability R is calculated as:
-$\text{R}=206.835-1.015\times\left(\frac{TW}{TSE}\right)-84.6\times\left(\frac{TSY}{TW}\right)$ (1)
-where, $TW$ is the total number of words in the given text, $TSE$ is the total number of sentences, and $TSY$ is the total number of syllables.
-$TW/TSE$ calculates the average sentence length: a higher average sentence length makes the text harder to read, so it subtracts from the score.
-$TSY/TW$ calculates the average syllables per word: more syllables per word reduce the score because longer words tend to make the text more difficult to read.
-Moreover, the compression rate C is calculated as:
-$\text{C}=\text{length}(Summary)/\text{length}(Text)$ (2)
-which is used to control the length of the generated summary.
-3.2.3 Test Process on Summarization Dataset
-According to the bottom of Figure 2, during the test process, LLM would first use $Exp_{QA}$ learned from the training process to generated auxiliary question-answer pairs, and bring them with the updated $Exp_{Sum}$ to generate the final summary.
-Prompts in test process are given in Appendix A.4.
-3.2.4 Test Process on QA Dataset
-In particular, as known from Figure 2.
-We design QG, a more challenging text generation strategy for LLM, to replace the text understanding strategy of predicting answers for the questions from the given QA dataset during the training process, in order to better fit LLM to the summarization task.
-A by-product of this approach is that we can even run the QA dataset in test process to generate summaries for its own texts based on its QA pairs, and adopt appropriate metrics for effectiveness evaluation, as shown in Figure 3.
-Prompts in test process for QA dataset are given in Appendix A.5.
-4 Experiments
-This section presents experimental details on four datasets, including two narrative dataset and two news datasets.
-We compare the quality of the summaries generated by the baseline models and our model using the generation algorithm from different aspects and perform an ablation study to analyze the influence of different components of T3.
-4.1 Settings
-4.1.1 Datasets
-Four datasets are used, including two summarization ones (BBC summary and NarraSum), and two QA ones (NLQuAD and FairytaleQA):
-BBC Summary includes 2,225 BBC news from 2004-2005, used for categorization and summary research Greene and Cunningham (2006).
-NarraSum has 122K narrative synopses from movies and TV episodes, aimed at improving summarization research Zhao et al. (2022).
-NLQuAD contains 31,000 questions and answers derived from 13,000 BBC news, for non-factoid long QA, focusing on multi-sentence descriptive answers to enhance document-level language models and challenges models with long context and complex language understanding Soleimani et al.
-(2021).
-FairytaleQA includes 278 children’s stories from Project Gutenberg, with questions and answers created by educational experts, aimed at assessing and training narrative comprehension for students (K-8) and models Xu et al.
-(2022).
-In experiments, 386 long articles from BBC summary, 260 long stories from FairytaleQA, 100 long stories from NarraSum, 501 long articles from NLQuAD are randomly picked up for evaluation.
-Each selected texts has a median word count exceeding 1,000.
-4.1.2 Evaluation Metrics
-To make more comprehensive evaluation, especially focusing on generated text quality and hallucinations alleviation, a range of classical and innovative metrics are used here:
-ROUGE compares the overlap of n-grams, word sequences, and word pairs between the generated summary and reference summaries Lin (2004).
-BLEU compares the overlap of n-grams between the generated text and reference translations, with higher scores indicating closer alignment with human translations and thus better quality Papineni et al.
-(2002).
-FactSumm FactSum uses named-entity recognition and relation extraction to extract facts from the source text and the generated summary to evaluate fact consistency by comparing extracted facts represented as triples (subject, relation, object) Zhang et al.
-(2021).
-ChatGPT evaluates factual inconsistency in summaries, outperforming previous methods with fewer hallucinations Luo et al. (2023).
-4.1.3 Baseline Models
-Seven advanced LLMs from three most representative series are selected as the baselines.
-All of them are accessed through Application Programming Interface (API), which not only standardizes the experiment for reproducibility and consistency, but also saves computational resources.
-GPT Series are developed by OpenAI and show strong language comprehension and generation capabilities for various natural language processing tasks Achiam et al.
-(2023).
-Here, GPT-3.5-turbo, GPT-4 Turbo, and the newest lauched GPT-4o are adopted here.
-Claude Series introduce advanced multimodal capabilities developed by Anthropic Anthropic (2023b).
-Claude 3.5 Sonnet is the newest one of them, showing strong performance in graduate-level and text-based reasoning Anthropic (2023a).
-Claude 3.5 Sonnet and Claude 3.0 Sonnet are chosen here.
-Gemini Series surpasses Google’s capabilities in understanding images, audio, video, and text Team et al.
-(2023).
-Currently, Gemini-1.5-pro is the newest and best-performing one of them Reid et al.
-(2024).
-Both Gemini-1.5-pro and Gemini-1.0-pro are used here.
-Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
-BBC summary
-w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
-GPT-4o 0.4019 | 0.4171 0.1814 | 0.2145 0.3697 | 0.3944 0.1086 | 0.1094 0.6364 | 0.7012
-GPT-4-turbo 0.3733 | 0.3683 0.1476 | 0.1414 0.3373 | 0.3327 0.0851 | 0.0769 0.6091 | 0.6901
-GPT-3.5-turbo 0.3936 | 0.3788 0.1775 | 0.1695 0.3604 | 0.3478 0.1050 | 0.0858 0.5752 | 0.4780
-Gemini-1.5-pro 0.3362 | 0.3510 0.1148 | 0.1258 0.3036 | 0.3169 0.0505 | 0.0647 0.2588 | 0.6049
-Gemini-1.0-pro 0.3490 | 0.3358 0.1280 | 0.1392 0.3175 | 0.3117 0.0599 | 0.0722 0.3576 | 0.5561
-Claude-3.5-sonnet 0.3908 | 0.4354 0.1659 | 0.2113 0.3588 | 0.4036 0.0970 | 0.1300 0.6343 | 0.6583
-Claude-3.0-sonnet 0.4123 | 0.4090 0.1826 | 0.1808 0.3778 | 0.3739 0.1108 | 0.1074 0.6002 | 0.5043
-NarraSum
-w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
-GPT-4o 0.3114 | 0.2886 0.0731 | 0.0858 0.2844 | 0.2566 0.0268 | 0.0255 0.9775 | 0.9969
-GPT-4-turbo 0.2914 | 0.2727 0.0772 | 0.0742 0.2544 | 0.2445 0.0336 | 0.0360 0.9775 | 0.9744
-GPT-3.5-turbo 0.2860 | 0.2528 0.0752 | 0.0588 0.2650 | 0.2300 0.0279 | 0.0215 0.9520 | 0.9591
-Gemini-1.5-pro 0.2599| 0.2695 0.0521 | 0.0539 0.2500 | 0.2532 0.0175 | 0.0199 0.9326 | 0.9357
-Gemini-1.0-pro 0.2897 | 0.2825 0.0626 | 0.0574 0.2545 | 0.2636 0.0165 | 0.0209 0.9285 | 0.9122
-Claude-3.5-sonnet 0.2919 | 0.2978 0.0717 | 0.0779 0.2673 | 0.2723 0.0310 | 0.0323 0.9704 | 0.9814
-Claude-3.0-sonnet 0.2990 | 0.2587 0.0675 | 0.0547 0.2645 | 0.2251 0.0274 | 0.0205 0.9469 | 0.9561
-NLQuAD
-w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
-GPT-4o 0.2437 | 0.3100 0.0848 | 0.1303 0.2271 | 0.2927 0.0033 | 0.0142 0.5981 | 0.6395
-GPT-4-turbo 0.2419 | 0.2366 0.0739 | 0.0736 0.2216 | 0.2162 0.0044 | 0.0036 0.5333 | 0.4916
-GPT-3.5-turbo 0.2150 | 0.2221 0.0775 | 0.0830 0.2007 | 0.2069 0.0027 | 0.0027 0.5530 | 0.4464
-Gemini-1.5-pro 0.2299 | 0.2419 0.0697 | 0.0797 0.2134 | 0.2244 0.0025 | 0.0047 0.5085 | 0.6509
-Gemini-1.0-pro 0.2626 | 0.2811 0.0929 | 0.1105 0.2455 | 0.2648 0.0054 | 0.0125 0.6457 | 0.7164
-Claude-3.5-sonnet 0.3203 | 0.4017 0.1280 | 0.1874 0.3006 | 0.3792 0.0235 | 0.0500 0.6571 | 0.8276
-Claude-3.0-sonnet 0.3235 | 0.3372 0.1270 | 0.1327 0.3016 | 0.3122 0.0197 | 0.0237 0.5889 | 0.6169
-FairytaleQA
-w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
-GPT-4o 0.1334 | 0.1486 0.0315 | 0.0372 0.1234 | 0.1378 0.0004 | 0.0007 0.9855 | 0.9892
-GPT-4-turbo 0.1468 | 0.1414 0.0351 | 0.0336 0.1345 | 0.1284 0.0009 | 0.0010 0.9744 | 0.9767
-GPT-3.5-turbo 0.1222 | 0.1223 0.0298 | 0.0307 0.1131 | 0.1074 0.0003 | 0.0004 0.9982 | 0.9737
-Gemini-1.5-pro 0.1540 | 0.1756 0.0380 | 0.0434 0.1430 | 0.1546 0.0010 | 0.0013 0.9723 | 0.9853
-Gemini-1.0-pro 0.1414 | 0.1343 0.0325 | 0.0243 0.1422 | 0.1339 0.0003 | 0.0014 0.9784 | 0.9424
-Claude-3.5-sonnet 0.1597 | 0.2096 0.0373 | 0.0607 0.1494 | 0.1952 0.0017 | 0.0047 0.9956 | 0.9997
-Claude-3.0-sonnet 0.1720 | 0.1891 0.0408 | 0.0490 0.1581 | 0.1735 0.0017 | 0.0033 0.9747 | 0.9892
-Table 1: Experimental Results on BBC summary, NarraSum, NLQuAD and FairytaleQA datasets where each LLM without (left col) or with (right col) T3.
-4.2 Implementation Details
-The comparative experiment are designed and executed to validate both the feasibility and effectiveness of the proposed T3, by investigating the generated summarization results for all four datasets from all used LLMs with/without using T3.
-The baseline summary of each text for each LLM is generated by using original text and the following prompt: "You are one helpful text assistant.
-Based on the input text, provide a summary in proper length".
-Correspondently, each LLM with T3 is guided by the experience and prompts derived from training process.
-For BBC summary and NarraSum in test process, both the original texts from them, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
-For FairytaleQA and NLQuAD in test process, 10 texts and their corresponding QA pairs from each dataset are selected for training, while the remaining texts are used for testing.
-Both the original texts from the datasets, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
-The higher scores from the used metrics are, indicating the better generation of summaries.
-It should be noted that the metric Factscore was calculated by GPT-4o for NarraSum and FairytaleQA, while calculated by FactSumm for BBC summary and NLQuAD, because it cannot extract facts from the story source text in practice so that does not perform well on narratives.
-For BBC summary and NarraSum, the provided summary was used as the reference, while for FairytaleQA and NLQuAD datasets, the original text was used as the reference.
-As mentioned in Part 3.2.4, T3 can also deal with QA datasets for summarization task.
-When the dataset does not provide reference summaries, the generated summary can be directly compared to the original text through certain automated metrics tools (e.g., ROUGE) Louis and Nenkova (2009).
-According to Saggion and Poibeau (2013), the overlap of the generated summaries with the original text is important in assessing information coverage and the validity.
-Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Readbility Factscore
-Claude-3.5-sonnet BBC summary
-Full T3 0.4354 0.2113 0.4036 0.1300 54.71 0.6583
-w/o summary experience 0.4026 0.1683 0.3599 0.1331 55.08 0.5832
-w/o qa experience 0.4155 0.2072 0.4005 0.3190 55.64 0.5857
-GPT-4o NLQuAD
-Full T3 0.3100 0.1303 0.2927 0.0142 44.53 0.6395
-w/o summary experience 0.3628 0.1529 0.3402 0.0309 42.57 0.7785
-w/o qa experience 0.4116 0.2029 0.3899 0.0531 45.77 0.6523
-Table 2: Ablation study results on BBC summary and NLQuAD datasets.
-For models from GPT or Claude series, default values for temperature and other related hyper-parameters are applied.
-For Gemini models, we set ’HARASSMENT’, ’HATE_SPEECH’, ’SEXUALLY_EXPLICIT’, ’DANGEROUS’ in safety setting into ’block_none’.
-4.3 Results
-Table 1 illustrates the performance of all used models with (right "T3" column) and without (left "w/o" column) T3 on all adopted datasets.
-The better one of the two values from each row of the two columns is marked in bold.
-Especially, for each row where all results from columns with T3 are better than those columns without T3, the row is colored blue.
-From this table, we observe that:
-(1) LLMs using T3 outperforms themselves without T3 under most metrics of most datasets, which demonstates the feasibility and effectiveness of T3.
-Specially, the average improvement rates of GPT-4o, Gemini-1.5-pro, and Claude-3.5-sonnet using T3 compared to the baselines without T3 across different tasks are calculated and shown that these LLMs with T3 nearly improves the ROUGE score by 14%, the BLEU score by 35%, and the Factscore score by 16%.
-(2) Of all the models in the four series, it is almost always the newest ones with T3 that achieve the most significant results comparing with itself without T3.
-For example, GPT-4o is 4% and 5% higher on Rouge-1 metrics compared to GPT-4 and GPT-3.5 on BBC Summary, respectively.
-These results could be due to the improvement of capabilities for latest LLMs on handling longer texts and more comprehensive tasks through understanding more complicated prompts and locating better details and events within lengthy texts Anthropic (2023a); Islam and Moushi (2024); Reid et al.
-(2024).
-(3) Four used datasets are divided into news- and narrative-styles.
-For most models with/without T3, the generated summaries on narrative-style datasets NarraSum seems not as good as that of news-style ones BBC Summary.
-According to Genette (1980); Van Dijk (2013); Barthes (2016), the structure of news-style texts is compared with that of narrative-style texts, highlighting how news is easier to be summarized because of its concise layout and fixed pattern, while the narrative texts is more complexity, including chronological order, shifts in point of view, and narrative hierarchy, so that increasing the difficulty of summarization.
-(4) The improvement of LLMs with T3 seems more significant on NLQuAD and FairytaleQA, i.e., QA task-oriented datasets, which might be benefited from using the prior information of questions and answers pairs to enhance the contextual analyzing capability of T3.
-Moreover, incorporating QA information implies to improve summarization quality by enhancing text comprehension, focusing on key information, increasing coverage, ensuring coherence, and aligning summaries with user needs Fabbri et al.
-(2022); Siledar et al.
-(2024).
-4.4 Ablation Study
-We further analyze the influence of different components of T3.
-Table 2 shows the experimental results for Claude-3.5-sonnet after removing summary experience and QA generation experience on the BBC summary dataset, and the results for GPT-4o after removing summary experience and QA generation experience on the NLQuAD dataset.
-For models with less reasoning ability, longer inputs and complex instructions of T3 might diminish their performance, while for more advanced models, each component of T3 could be utilized to enhance their overall effectiveness.
-5 Conclusion
-A novel T3 framework is proposed to address the challenge of long text summarization, using QA datasets to iteratively train baseline model at first, then to deal with summarization in a zero-shot way.
-Experiments on four datasets among seven representative baselines under five evaluation metrics validate the significant effectiveness of T3.
-Since T3 is model-agnostic and theoritically could be used for any suitable assistant task and target task pair, the future work will aim to explore the potential of it in a broader range of task combinations.
-Limitations
-When using the proposed T3 framework in real scenarios, the following limitations that need to be further optimized:
-Limited Baseline Models Due to the costs and time constraints of API, we only test T3 on models from GPT, Claude, and Gemini series.
-More open-source models (such as LLaMA) and their private deployment situation will be considered in the future.
-Limited Tasks Few empirical guidelines for the selection of assistant tasks are given in this paper.
-In the future, the matching rules of assistant tasks and target tasks will be comprehensively explored in various task combinations.
-Limited Evaluation Metrics As to evaluate generation results, BLEU and ROUGE focus on word overlap and neglect fluency, semantics, and factual accuracy, while FactSumm aims to consistency but struggles with abstract concepts.
-For summarization task, more suitable evaluation metrics should be discussed.
-Ethics Statement
-Data Availability and Safety The training and test data used for summarization and QA in this paper are publicly available.
-While the original datasets were filtered, some content may still include sensitive material, such as reports on violent crimes or incidents.
-Furthermore, when utilizing the Gemini model service, the safety mode is set to ’block_none,’ which may result in outputs containing violent or explicit content.
-Usage of Large Language Models Models like GPT, Gemini, and Claude are utilized to generate QA results, refine prompt descriptions, and produce text summaries for input texts in summarization tasks.
-The generated text is used exclusively for experimental and analytical purposes, and the findings are presented in the relevant sections.
-Appendix A Appendix
-A.1 Training Steps of T3 for Summarization
+Furthermore, the detailed information of the training process is given in Algorithm 1, including the following key steps:
 Algorithm 1 T3 Training for Summarization
@@
 15:  end for
+Step 1 Updating $Exp_{QA}$: At first, both summary and QA generation experiences $Exp_{Sum}$ and $Exp_{QA}$ are initialized as $null$.
+Then, LLM generates new QA pairs $QA_{i}^{\prime}$ for each text $D_{i}$ and updates $Exp_{QA}$ according to several rules concluded by itself through comparing $QA_{i}\textquoteright$ ’ with the ground truth $QA_{i}$.
+The prompt $P_{QA}$ is predefined to control all steps mentioned in this part.
+Step 2 Updating $Exp_{Sum}$: In this part, LLM follows the predefined prompts $P_{Sum}$ and try to summarize the text $D_{i}$.
+As an prior-knowledge that might bring extra contextual information, both $QA_{i}$ and $QA_{i}\textquoteright$ ’ are also taken into account as the inputs.
+After the current summary $Sum_{i}^{\prime}$ is generated, three thresholds are used to investigate whether the generated content is well enough.
+LLM would repeat the summary generation step and update its experience $Exp_{Sum}$ every iteration, till the stop conditions are fulfilled.
+Step 3 Thresholds Used in Iteration: The iterative process mentioned above continues until the generated summary $Sum_{i}^{\prime}$ simultaneously meets the required thresholds for similarity, readability Flesch (1948), and compression rate (or the number of iterations reach the preset maximum iteration times $K$).
+In details, readability R is calculated as:
+$\text{R}=206.835-1.015\times\left(\frac{TW}{TSE}\right)-84.6\times\left(\frac{TSY}{TW}\right)$ (1)
+where, $TW$ is the total number of words in the given text, $TSE$ is the total number of sentences, and $TSY$ is the total number of syllables.
+$TW/TSE$ calculates the average sentence length: a higher average sentence length makes the text harder to read, so it subtracts from the score.
+$TSY/TW$ calculates the average syllables per word: more syllables per word reduce the score because longer words tend to make the text more difficult to read.
+Moreover, the compression rate C is calculated as:
+$\text{C}=\text{length}(Summary)/\text{length}(Text)$ (2)
+which is used to control the length of the generated summary.
+3.3 Test Process on Summarization Dataset
+According to the bottom of Figure 1, during the test process, LLM would first use $Exp_{QA}$ learned from the training process to generated auxiliary question-answer pairs, and bring them with the updated $Exp_{Sum}$ to generate the final summary.
+Prompts in test process are given in Appendix A.4 and a step-by-step use case is shown in Appendix A.10.3.
+3.4 Test Process on QA Dataset
+In particular, as known from Figure 1.
+We design QG, a more challenging text generation strategy for LLM, to replace the text understanding strategy of predicting answers for the questions from the given QA dataset during the training process, in order to better fit LLM to the summarization task.
+A by-product of this approach is that we can even run the QA dataset in test process to generate summaries for its own texts based on its QA pairs, and adopt appropriate metrics for effectiveness evaluation, as shown in Figure 2.
+Prompts in test process for QA dataset are given in Appendix A.5.
+Figure 2: Test process of T3 for summarization task on QA dataset.
+3.5 T3 General Workflow
+The keypoint of T3 could be concluded that: in the whole training process, both the assistant task and the target task share the samples of the assistant task.
+At first, the used LLM  [^1] GPT-4o is recommended, while any other large models with sufficient parameter capacity to possess emergence could be considered.
+acquires experience in handling the assistant task through the labeled data, which means a deeper understanding of the samples.
+Subsequently, LLM iteratively tries to accomplish the target task on these same samples and gain the experience for the target task by leveraging the previous learned contextual experiences, as well as being modulated by the predefined conditions.
+Finally, the model applies the experiences required for both the assistant and target tasks during the test process.
+Accordingly, a general workflow for task-agnostic T3 is shown in Figure 3 from Appendix A.1.
+4 Experiments
+This section presents experimental details on four datasets, including two narrative dataset and two news datasets.
+We compare the quality of the summaries generated by the baseline models and our model using the generation algorithm from different aspects and perform an ablation study to analyze the influence of different components of T3.
+4.1 Settings
+4.1.1 Datasets
+Four datasets are used, including two summarization ones (BBC summary and NarraSum), and two QA ones (NLQuAD and FairytaleQA):
+BBC Summary includes 2,225 BBC news from 2004-2005, used for categorization and summary research Greene and Cunningham (2006).
+NarraSum has 122K narrative synopses from movies and TV episodes, aimed at improving summarization research Zhao et al. (2022).
+NLQuAD contains 31,000 questions and answers derived from 13,000 BBC news, for non-factoid long QA, focusing on multi-sentence descriptive answers to enhance document-level language models and challenges models with long context and complex language understanding Soleimani et al.
+(2021).
+FairytaleQA includes 278 children’s stories from Project Gutenberg, with questions and answers created by educational experts, aimed at assessing and training narrative comprehension for students (K-8) and models Xu et al.
+(2022).
+In experiments, 386 long articles from BBC summary, 260 long stories from FairytaleQA, 100 long stories from NarraSum, 501 long articles from NLQuAD are randomly picked up for evaluation.
+Each selected texts has a median word count exceeding 1,000.
+4.1.2 Evaluation Metrics
+To make more comprehensive evaluation, especially focusing on generated text quality and hallucinations alleviation, a range of classical and innovative metrics are used here:
+ROUGE compares the overlap of n-grams, word sequences, and word pairs between the generated summary and reference summaries Lin (2004).
+BLEU compares the overlap of n-grams between the generated text and reference translations, with higher scores indicating closer alignment with human translations and thus better quality Papineni et al.
+(2002).
+Factscore We use FactSumm and ChatGPT for Factscore calculation.
+FactSumm uses named-entity recognition and relation extraction to extract facts from the source text and the generated summary to evaluate fact consistency by comparing extracted facts represented as triples (subject, relation, object) Zhang et al.
+(2021).
+ChatGPT evaluates factual inconsistency in summaries, outperforming previous methods with fewer hallucinations Luo et al.
+(2023).
+4.1.3 Baseline Models
+Seven advanced LLMs from three most representative series are selected as the baselines.
+All of them are accessed through Application Programming Interface (API), which not only standardizes the experiment for reproducibility and consistency, but also saves computational resources.
+GPT Series are developed by OpenAI and show strong language comprehension and generation capabilities for various natural language processing tasks Achiam et al.
+(2023).
+Here, GPT-3.5-turbo, GPT-4 Turbo, and the newest lauched GPT-4o are adopted here.
+Claude Series introduce advanced multimodal capabilities developed by Anthropic Anthropic (2023b).
+Claude 3.5 Sonnet is the newest one of them, showing strong performance in graduate-level and text-based reasoning Anthropic (2023a).
+Claude 3.5 Sonnet and Claude 3.0 Sonnet are chosen here.
+Gemini Series surpasses Google’s capabilities in understanding images, audio, video, and text Team et al.
+(2023).
+Currently, Gemini-1.5-pro is the newest and best-performing one of them Reid et al.
+(2024).
+Both Gemini-1.5-pro and Gemini-1.0-pro are used here.
+Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
+BBC summary
+w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
+GPT-4o* 40.19 | 41.71 18.14 | 21.45 36.97 | 39.44 10.86 | 10.94 63.64 | 70.12
+GPT-4-turbo 37.33 | 36.83 14.76 | 14.14 33.73 | 33.27 8.51 | 7.69 60.91 | 69.01
+GPT-3.5-turbo* 39.36 | 37.88 17.75 | 16.95 36.04 | 34.78 10.50 | 8.58 57.52 | 47.80
+Gemini-1.5-pro* 33.62 | 35.10 11.48 | 12.58 30.36 | 31.69 5.05 | 6.47 25.88 | 60.49
+Gemini-1.0-pro* 34.90 | 33.58 12.80 | 13.92 31.75 | 31.17 5.99 | 7.22 35.76 | 55.61
+Claude-3.5-sonnet* 39.08 | 43.54 16.59 | 21.13 35.88 | 40.36 9.70 | 13.00 63.43 | 65.83
+Claude-3.0-sonnet 41.23 | 40.90 18.26 | 18.08 37.78 | 37.39 11.08 | 10.74 60.02 | 50.43
+Average 37.96 | 38.51 15.68 | 16.89 34.64 | 35.44 8.81 | 9.23 52.45 | 59.90
+NarraSum
+w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
+GPT-4o* 31.14 | 28.86 7.31 | 8.58 28.44 | 25.66 2.68 | 2.55 97.75 | 99.69
+GPT-4-turbo* 29.14 | 27.27 7.72 | 7.42 25.44 | 24.45 3.36 | 3.60 97.75 | 97.44
+GPT-3.5-turbo* 28.60 | 25.28 7.52 | 5.88 26.50 | 23.00 2.79 | 2.15 95.20 | 95.91
+Gemini-1.5-pro* 25.99 | 26.95 5.21 | 5.39 25.00 | 25.32 1.75 | 1.99 93.26 | 93.57
+Gemini-1.0-pro 28.97 | 28.25 6.26 | 5.74 25.45 | 26.36 1.65 | 2.09 92.85 | 91.22
+Claude-3.5-sonnet 29.19 | 29.78 7.17 | 7.79 26.73 | 27.23 3.10 | 3.23 97.04 | 98.14
+Claude-3.0-sonnet* 29.90 | 25.87 6.75 | 5.47 26.45 | 22.51 2.74 | 2.05 94.69 | 95.61
+Average 28.99 | 27.47 6.85 | 6.61 26.29 | 24.93 2.58 | 2.52 95.51 | 95.94
+NLQuAD
+w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
+GPT-4o* 24.37 | 31.00 8.48 | 13.03 22.71 | 29.27 0.33 | 1.42 59.81 | 63.95
+GPT-4-turbo* 24.19 | 23.66 7.39 | 7.36 22.16 | 21.62 0.44 | 0.36 53.33 | 49.16
+GPT-3.5-turbo* 21.50 | 22.21 7.75 | 8.30 20.07 | 20.69 0.27 | 0.27 55.30 | 44.64
+Gemini-1.5-pro* 22.99 | 24.19 6.97 | 7.97 21.34 | 22.44 0.25 | 0.47 50.85 | 65.09
+Gemini-1.0-pro* 26.26 | 28.11 9.29 | 11.05 24.55 | 26.48 0.54 | 1.25 64.57 | 71.64
+Claude-3.5-sonnet* 32.03 | 40.17 12.80 | 18.74 30.06 | 37.92 2.35 | 5.00 65.71 | 82.76
+Claude-3.0-sonnet* 32.35 | 33.72 12.70 | 13.27 30.16 | 31.22 1.97 | 2.37 58.89 | 61.69
+Average 26.24 | 29.01 9.34 | 11.39 24.44 | 27.09 0.88 | 1.59 58.35 | 62.70
+FairytaleQA
+w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3 w/o   |    T3
+GPT-4o* 13.34 | 14.86 3.15 | 3.72 12.34 | 13.78 0.04 | 0.07 98.55 | 98.92
+GPT-4-turbo* 14.68 | 14.14 3.51 | 3.36 13.45 | 12.84 0.09 | 0.10 97.44 | 97.67
+GPT-3.5-turbo 12.22 | 12.23 2.98 | 3.07 11.31 | 10.74 0.03 | 0.04 99.82 | 97.37
+Gemini-1.5-pro* 15.40 | 17.56 3.80 | 4.34 14.30 | 15.46 0.10 | 0.13 97.23 | 98.53
+Gemini-1.0-pro* 14.14 | 13.43 3.25 | 2.43 14.22 | 13.39 0.03 | 0.14 97.84 | 94.24
+Claude-3.5-sonnet* 15.97 | 20.96 3.73 | 6.07 14.94 | 19.52 0.17 | 0.47 99.56 | 99.97
+Claude-3.0-sonnet* 17.20 | 18.91 4.08 | 4.90 15.81 | 17.35 0.17 | 0.33 97.47 | 98.92
+Average 14.71 | 16.01 3.50 | 3.98 13.77 | 14.73 0.09 | 0.18 98.27 | 97.95
+Table 1: Experimental Results on BBC summary, NarraSum, NLQuAD and FairytaleQA datasets where each LLM without (left col) or with (right col) T3.
+4.2 Implementation Details
+The comparative experiment are designed and executed to validate both the feasibility and effectiveness of the proposed T3, by investigating the generated summarization results for all four datasets from all used LLMs with/without using T3.
+The baseline summary of each text for each LLM is generated by using original text and the following prompt: "You are one helpful text assistant.
+Based on the input text, provide a summary in proper length.
+Aim to maximize Performance scores".
+[^2] We add this to ensure fairness, since T3 use these words in its training process, according to Appendix A.3.
+Correspondently, each LLM with T3 is guided by the experience and prompts derived from training process.
+For BBC summary and NarraSum in test process, both the original texts from them, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
+For FairytaleQA and NLQuAD in test process, 10 texts and their corresponding QA pairs from each dataset are selected for training, while the remaining texts are used for testing.
+Both the original texts from the datasets, the QA generation experience $Exp_{QA}$ and the summary experience $Exp_{Sum}$ are contained into the input prompt, whose detailed prompts are given in Appendix A.6, A.7, A.8 and A.9 for News-style Datasets and Narratives-style Datasets respectively.
+The higher scores from the used metrics are, indicating the better generation of summaries.
+It should be noted that the metric Factscore was calculated by GPT-4o for NarraSum and FairytaleQA, while calculated by FactSumm for BBC summary and NLQuAD, because it cannot extract facts from the story source text in practice so that does not perform well on narratives.
+For BBC summary and NarraSum, the provided summary was used as the reference, while for FairytaleQA and NLQuAD datasets, the original text was used as the reference.
+As mentioned in Part 3.4, T3 can also deal with QA datasets for summarization task.
+When the dataset does not provide reference summaries, the generated summary can be directly compared to the original text through certain automated metrics tools (e.g., ROUGE) Louis and Nenkova (2009).
+According to Saggion and Poibeau (2013), the overlap of the generated summaries with the original text is important in assessing information coverage and the validity.
+For models from GPT or Claude series, default values for temperature and other related hyper-parameters are applied.
+For Gemini models, we set ’HARASSMENT’, ’HATE_SPEECH’, ’SEXUALLY_EXPLICIT’, ’DANGEROUS’ in safety setting into ’block_none’.
+4.3 Results
+Table 1 illustrates the performance of all used models with (right "T3" column) and without (left "w/o" column) T3 on all adopted datasets.
+The better one of the two values from each row of the two columns is marked in bold.
+Especially, for each row where all results from columns with T3 are better than those columns without T3, the row is colored blue.
+Models corresponding to results with significant differences between w/o and T3 are superscripted with *, where the detailed calculation process is given in Appendix A.12.
+From this table, we observe that:
+(1) LLMs using T3 outperforms themselves without T3 under most metrics of most datasets, which demonstates the feasibility and effectiveness of T3.
+Specially, the average improvement rates of GPT-4o, Gemini-1.5-pro, and Claude-3.5-sonnet using T3 compared to the baselines without T3 across different tasks are calculated and shown that these LLMs with T3 nearly improves the ROUGE score by 14%, the BLEU score by 35%, and the Factscore score by 16%.
+Most of the comparison groups (82.14%) showed significant differences with significance_level=0.05 and null_hypothesis that there is no significant difference between the means of the two groups.
+(2) Of all the models in the four series, it is almost always the newest ones with T3 that achieve the most significant results comparing with itself without T3.
+For example, GPT-4o is 4% and 5% higher on Rouge-1 metrics compared to GPT-4 and GPT-3.5 on BBC Summary, respectively.
+These results could be due to the improvement of capabilities for latest LLMs on handling longer texts and more comprehensive tasks through understanding more complicated prompts and locating better details and events within lengthy texts Anthropic (2023a); Islam and Moushi (2024); Reid et al.
+(2024).
+(3) Four used datasets are divided into news- and narrative-styles.
+For most models with/without T3, the generated summaries on narrative-style datasets NarraSum seems not as good as that of news-style ones BBC Summary.
+According to Genette (1980); Van Dijk (2013); Barthes (2016), the structure of news-style texts is compared with that of narrative-style texts, highlighting how news is easier to be summarized because of its concise layout and fixed pattern, while the narrative texts is more complexity, including chronological order, shifts in point of view, and narrative hierarchy, so that increasing the difficulty of summarization.
+(4) The improvement of LLMs with T3 seems more significant on NLQuAD and FairytaleQA, i.e., QA task-oriented datasets, which might be benefited from using the prior information of questions and answers pairs to enhance the contextual analyzing capability of T3.
+Moreover, incorporating QA information implies to improve summarization quality by enhancing text comprehension, focusing on key information, increasing coverage, ensuring coherence, and aligning summaries with user needs Fabbri et al.
+(2022); Siledar et al.
+(2024).
+4.4 Ablation Study
+We further analyze the influence of different components of T3.
+Table LABEL:tab:Table3 in Appendix A.11 shows the experimental results for all used models after removing the summary experience component $Exp_{Sum}$ or the QA generation experience component $Exp_{QA}$, on all used datasets.
+It could be found that: (1) For models with less reasoning ability, longer inputs and complex instructions might diminish their performance mainly in their $Exp_{Sum}$ component than their $Exp_{QA}$ component, while for more advanced models, both the two components could be utilized to enhance their overall effectiveness.
+(2) Since $Exp_{Sum}$ and $Exp_{QA}$ is run for the target task and the assistant task, respectively, it might be demonstrated the importance of the assistant task in the whole T3 workflow.
+5 Conclusion
+A novel T3 framework is proposed to address the challenge of long text summarization, using QA datasets to iteratively train baseline model at first, then to deal with summarization in a zero-shot way.
+Experiments on four datasets among seven representative baselines under five evaluation metrics validate the significant effectiveness of T3.
+Since T3 is model-agnostic and theoritically could be used for any suitable assistant task and target task pair, the future work will aim to explore the potential of it in a broader range of task combinations.
+Limitations
+When using the proposed T3 framework in real scenarios, the following limitations that need to be further optimized:
+Limited Baseline Models Due to the costs and time constraints of API, we only test T3 on models from GPT, Claude, and Gemini series.
+More open-source models (such as LLaMA) and their private deployment situation will be considered in the future.
+Limited Tasks Few empirical guidelines for the selection of assistant tasks are given in this paper.
+In the future, the matching rules of assistant tasks and target tasks will be comprehensively explored in various task combinations.
+Limited Evaluation Metrics As to evaluate generation results, BLEU and ROUGE focus on word overlap and neglect fluency, semantics, and factual accuracy, while FactSumm aims to consistency but struggles with abstract concepts.
+For summarization task, more suitable evaluation metrics should be discussed.
+Ethics Statement
+Data Availability and Safety The training and test data used for summarization and QA in this paper are publicly available.
+While the original datasets were filtered, some content may still include sensitive material, such as reports on violent crimes or incidents.
+Furthermore, when utilizing the Gemini model service, the safety mode is set to ’block_none,’ which may result in outputs containing violent or explicit content.
+Usage of Large Language Models Models like GPT, Gemini, and Claude are utilized to generate QA results, refine prompt descriptions, and produce text summaries for input texts in summarization tasks.
+The generated text is used exclusively for experimental and analytical purposes, and the findings are presented in the relevant sections.
+Appendix A Appendix
+A.1 General workflow of task-agnostic T3
+The figure below shows the general task-agnostic workflow of T3, i.e., the overall workflow of T3 when both the assistant and target tasks are not identified.
+The workflow facilitates the matching of different assistant and target tasks in practice.
+Figure 3: General workflow of task-agnostic T3
 A.2 Prompt for QA Generation and Experience Updating in Training Process
-You are one helpful text assistant and you are good at learning new experience. First please help us to do the following:
+You are a helpful text assistant skilled at learning new experiences. First, please help us with the following:
+Based on the input text, provided questions, and QA pair generation experience, generate questions and corresponding answers that reflect all subjects and key points of the input article.
+The format of the generated QA pairs should match that of the provided question pairs.
+Update the learning experience for generating better QA pairs during the question generation process by comparing the generated questions to the original question pairs based on the original text or other methods.
+Consider all key points needed to create good QA pairs, and ensure the learning experience is general.
+DO NOT INCLUDE DETAILED EXAMPLES BUT PROVIDE RULES AS DETAILED AS POSSIBLE.
+DO NOT EXCEED 600 WORDS.
+The format of the output is as follows:
+{ "Generated_QA_pairs": { "1": "Question": "xxxx", "Answer": "xxxx" , "2": "Question": "xxxx", "Answer": "xxxx" }, "QA_generation_experience": "Answer point by point text in one line" }
+A.3 Prompt for Summarization Generation and Experience Updating in Training Process
+You are a helpful text assistant. Please help us with the following:
 1.
-Based on input text, provided questions and QA pair generation experience, generate questions and corresponding answers to original text that can reflect all the subjects and all the key points of input article.
-The format of the generated question pairs should be same as the provided question pairs.
+Based on the input text, provided questions, generated QA pairs, and summary generation experience, generate a summary that corresponds to the original text, reflecting all subjects and key points of the input article at an appropriate length.
 2.
-Update the learning experience of how to generate better question pairs during the question generation process by comparing the generated questions and original question pairs based on the original text or other methods.
-You need to consider all the key points to make a good QA pairs and the learning experience should be general.
-DO NOT INCLUDE DETAIL EXAMPLES BUT PROVIDE rules AS DETAILED AS POSSIBLE.DO NOT EXCEED 600 WORDS.
+Update the summary generation experience with insights on how to create a better summary by first using the provided questions, then the generated summary, and all provided questions (both generated and original), and finally incorporating what you’ve learned during question generation in the question generation experience.
+Here is the target for the generated summary:
+THE GENERATED SUMMARY SHOULD HAVE HIGHER PERFORMANCE SCORES COMPARED TO THE DIRECTLY GENERATED ONES.
+The summary should include essential terms, phrases, and sentences that closely mirror the original text.
+Aim to maximize the overlap with the reference summary by closely following the wording, sentence structure, and style of the original text.
+Ensure that the summary is concise but comprehensive, avoiding unnecessary details or deviations.
+Also, consider the fluency of the summary.
+Consider all key points to make a good summary, and make the learning experience general (e.g., avoid losing details in the story, maintain sentence coherence, and list some points to improve accuracy and fluency).
+Additionally, you should list other rules that might be important, with each point as detailed as possible.
+You may also consider methods to ensure a better summary, such as voting.
+Answer in one line and point by point. DO NOT EXCEED 600 WORDS.
 The format of the output is as follows:
-{ "Generated_QA_pairs":
-{ "1": { "Question": "xxxx", "Answer": "xxxx"},
-"2": { "Question": "xxxx", "Answer": "xxxx"},
-"QA_generation_experience": "Answer point by point text in one line"}
-A.3 Prompt for Summarization Generation and Experience Updating in Training Process
-You are one helpful text assistant. Please help us to do the following:
-1.
-Based on input text, provided questions, generated QA pairs and summary generation experience, generate summary corresponding answers to original text that can reflect all the subjects and all the key points of input article in proper length.
-2.
-Update the summary generation experience of how to generate better summary with the help of provided questions first by the generated summary and all provided questions (the generated and the original) and then by what you learn when generating questions in question generation experience.
-Here is the target for the generated summary:
-THE GENERATED SUMMARY SHOULD HAS HIGHER ROUGE AND BLEU SCORES ALIGNED WITH THE ORIGINAL TEXT THAN THE DIRECTED GENERATED ONES.
-The summary should include essential terms, phrases, and sentences that closely mirror the original text to ensure high precision and recall.
-Notice Aim to maximize the overlap with the reference summary by closely following the wording, sentence structure, and style of the original text.
-Ensure that the summary is concise but comprehensive, avoiding unnecessary details or deviations.
-You need also consider the fluency of summary.
-You need to consider all the key points to make a good summary and the learning experience should be general.(for example, like does not lose the detail in the story and pay attention to the connection of sentences, listing some point to improve the accuracy and the fluency of the story, YOU STILL NEED TO LIST OTHER rules THAT MAY IMPORTANT,Every point should as detailed as possible).
-You can also think some methods to ensure a better summary, such as voting.
-Answer in one line and point by point. DO NOT EXCEED 600 WORDS.’ ’The format of the output is as follows:
 { "Summary": "text in one line",
@@
 A.4 Prompt in Test Process on Summarization Dataset
-You are a helpful text assistant. Please help us to do the following:
-Based on the input text, with the help of QA generation rules and summary rules, here are the steps to do the summary:
+You are a helpful text assistant. Please help us with the following:
+Based on the input text, and with the help of QA generation rules and summary rules, here are the steps to create the summary:
 First, generate QA pairs based on the generation rules.
-Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
+Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text at the appropriate length.
 Article: [Article]
@@
 A.5 Prompt in Test Process on QA Dataset
-You are a helpful text assistant. Please help us to do the following:
-Based on the input text and the questions given, with the help of QA generation rules and summary rules, here are the steps to do the summary:
+You are a helpful text assistant. Please help us with the following:
+Based on the input text and the questions given, and with the help of QA generation rules and summary rules, here are the steps to create the summary:
 First, generate QA pairs based on the generation rules.
-Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text in proper length.
+Then, based on the input text, provided questions, generated QA pairs, and the summary generation experience, generate a summary of the original text at the appropriate length.
 Article: [Article]
@@
 A.6 Summary Experience for News-style Datasets
-1. Ensure all key points such as the formation, evolution, and current status of Boko Haram, and the contrasting views on security are included.
-2. Maintain original wording and sentence structure for precision and recall.
+1. Ensure all key points, such as the formation, evolution, and current status of Boko Haram, as well as contrasting views on security, are included.
+2. Maintain the original wording and sentence structure for precision and recall.
 3. Avoid unnecessary details, focusing on core information for conciseness.
@@
 6. Consider using a voting system to select the best summary from multiple drafts.
-7. Verify the summary reflects the original text accurately without speculative information.
-8. Address distinct key points or subjects in each question for comprehensive coverage.
-9. Compare generated questions with original ones to ensure all critical aspects are covered.
-10. Ensure answers are directly supported by the text, avoiding speculative or inferred information.
+7. Verify that the summary reflects the original text accurately without speculative information.
+8. Address distinct key points or subjects in the summary for comprehensive coverage.
+9. Compare generated summary with original text to ensure all critical aspects are covered.
+10. Ensure summary are directly supported by the text, avoiding speculative or inferred information.
 A.7 Summary Experience for Narratives-style Datasets
-1. Ensure all key points and characters from the text are covered to avoid missing crucial details,use Details to Enhance Emotional Impact such as Incorporate symbolic details.
+1. Ensure all key points and characters from the text are covered to avoid missing crucial details; use details to enhance emotional impact, such as incorporating symbolic details.
 2. Maintain the original sequence of events to reflect the narrative flow accurately.
 3. Use language and phrases closely mirroring the original text for higher precision and recall.
-4. Keep the summary concise but comprehensive, avoiding unnecessary details while ensuring all significant events are included.
+4. Keep the summary concise yet comprehensive, avoiding unnecessary details while ensuring all significant events are included.
 5. Focus on fluency, ensuring the summary reads smoothly and logically. Preserve the logical flow and connection between events for coherence.
@@
 9. Emphasize themes and symbolism: Don’t just recount the plot; also highlight the underlying themes or symbolic meanings.
-10. Aim to maximize ROUGE and BLEU scores by closely following the wording, sentence structure, and style of the original text.
+10. Closely follow the wording, sentence structure, and style of the original text.
 11.
@@
 This adds depth to the summary and helps convey the underlying messages or morals of the story.
-It’s important not just to recount events but also to capture the feelings and ideas that these events evoke.
+It’s important not just to recount events but also to capture the feelings and ideas these events evoke.
 12.
-Ensure that no important details are lost that contribute to the storyś moral or lesson.
+Ensure that no important details contributing to the story’s moral or lesson are lost.
 These details are often critical to understanding the overall message of the text, so they should be preserved to maintain the integrity of the original story.
-13. Consider using a voting method to select the best summary from multiple versions, such as generate 3 versions and find the best one.
+13. Consider using a voting method to select the best summary from multiple versions, such as generating three versions and choosing the best one.
 A.8 QA Generation Experience for News-style Datasets
@@
 4. Compare generated questions with original ones to ensure all critical aspects are covered, and questions are clear and concise.
-5. Answers should be directly supported by the text and avoid speculative or inferred information. This approach ensures comprehensive and accurate QA pairs.
+5. Answers should be directly supported by the text and should avoid speculative or inferred information. This approach ensures comprehensive and accurate QA pairs.
 A.9 QA Generation Experience for Narratives-style Datasets
-1.Ensure each question addresses a unique key point or detail to avoid redundancy. Cover all significant events, characters, and details for comprehensive coverage.
+1. Ensure each question addresses a unique key point or detail to avoid redundancy. Cover all significant events, characters, and details for comprehensive coverage.
 2. Keep questions clear and specific to avoid ambiguity and ensure they can be answered directly from the text.
 3. Include a mix of factual, inferential, and interpretative questions to engage different levels of understanding.
-4. Review the text multiple times to ensure no key points are missed and questions flow logically.
+4. Review the text multiple times to ensure no key points are missed and that questions flow logically.
 5. Pay attention to character motivations, actions, and consequences to generate deeper questions.
-6. Consider the context and sequence of events to create questions reflecting the narrativeś progression.
+6. Consider the context and sequence of events to create questions reflecting the narrative’s progression.
 7. Ensure answers are concise and directly derived from the text for accuracy.
 8. Compare generated questions with existing ones to identify gaps or overlaps and refine accordingly.
-9. Aim to create questions that encourage critical thinking and deeper engagement with the text.’
+9. Aim to create questions that encourage critical thinking and deeper engagement with the text.
+A.10 Step-by-step example
+This section provides a detailed use case description, which helps readers understand the operation flow of each step more intuitively by analyzing the application of T3 step by step.
+A.10.1 Original Text
+In this example, we apply one sample from the dataset NarraSum, a story containing more than 1700 words, as our summarization target.
+"The opening scene depicts Anita N̈eedyL̈esnicky (Amanda Seyfriend), a violent inmate in a correctional facility.
+She is notorious for being a ’kicker’, and is now reflecting during lockdown on past events.
+Needy is a nerdy, substantially unpopular, p̈lain Janeẗype student from Devil’s Kettle, Minnesota, who has been best friends with Jennifer Check (Megan Fox), a beautiful, popular cheerleader, since their childhood.
+…… The end credits show Low Shoulder checking into a hotel post-concert and candidly indulging in drugs and alcohol in their hotel room.
+Eventually, the band is savagely murdered by a blurred figure.
+We then see police crime scene photos of the band members’ corpses strewn around the hotel room, Nikolai having been stabbed in the chest with the sacrificial knife.
+The final image of the movie shows security camera footage of an evil-looking Needy walking down the hotel hallway past a gathering of female band groupies."
+A.10.2 Generate QA Pairs
+Using the sample and the QA generation experience in A.9 as input, the baseline model creates several question-answer pairs to capture key plot points and character details. Here are some examples:
+Question: "What is Needy’s situation at the beginning of the movie?"
+Answer: "She is an inmate in a correctional facility, known for her violent behavior as a ’kicker’."
+Question: "Who are the main characters and their relationship?"
+Answer: "Needy Lesnicky, an unpopular student, and Jennifer Check, a popular cheerleader, have been best friends since childhood.
+Question: "What happens to the band Low Shoulder at the end?"
+Answer: "The band is murdered by a figure, later revealed as Needy, after a concert while indulging in drugs and alcohol.
+A.10.3 Final Summary
+Using the generated QA pairs and the summary experience in A.7 as input, the baseline model generates a summary of the story as follows:
+The story revolves around Anita N̈eedyL̈esnicky, a violent inmate reflecting on past events.
+Needy, a nerdy student from Devil’s Kettle, Minnesota, was best friends with Jennifer Check, a popular cheerleader.
+One night, they attend a concert by indie rock band Low Shoulder at a local bar.
+The bar catches fire, and Jennifer leaves with the band despite Needy’s protests.
+Later, Jennifer appears at Needy’s house, injured and behaving strangely.
+The next day, Jennifer acts normal at school but soon begins killing and feeding on male students.
+Needy becomes suspicious and investigates, discovering that Low Shoulder had attempted to sacrifice Jennifer in a Satanic ritual for fame and success.
+However, because Jennifer wasn’t a virgin, the ritual backfired, turning her into a succubus.Needy tries to warn her boyfriend Chip about Jennifer’s true nature, but he doesn’t believe her.
+At the Spring Formal, Jennifer attacks Chip, leading to a confrontation between Needy and Jennifer.
+Chip dies in Needy’s arms after trying to save her.
+Consumed by revenge, Needy breaks into Jennifer’s house and kills her.
+Needy is institutionalized for Jennifer’s murder but develops supernatural abilities from being bitten during their fight.
+She escapes the institution and sets out to seek revenge on Low Shoulder.
+The story concludes with security footage showing Needy walking down a hotel hallway, implying she has murdered the band members.
+The narrative explores themes of friendship, betrayal, and the supernatural, while also critiquing the music industry and teenage social dynamics.
+It blends elements of horror, dark comedy, and coming-of-age drama to create a unique and compelling story.
+This final summary keeps the plot concise while covering essential character arcs, the supernatural element, and the revenge ending.
+As shown in this example, by using QA to capture plot intricacies, the proposed T3 framework ensures that the final summary is both comprehensive and aligned with the main themes of the original content.
+A.11 Ablation study
+Table 2: Ablation study results on BBC summary, NarraSum, NLQuAD and FairytaleQA
+Models Strategy ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
+GPT-3.5-turbo NLQuAD
+Full T3 22.21 8.30 20.69 0.27 47.80
+w/o summary 31.68 15.57 30.24 5.31 46.36
+w/o qa 26.01 9.59 23.91 0.83 45.35
+BBC Summary
+Full T3 37.88 16.95 34.78 8.58 47.80
+w/o summary 40.62 22.74 37.05 12.66 55.49
+w/o qa 36.00 15.07 32.39 8.14 46.08
+FairyTaleQA
+Full T3 12.23 3.07 10.74 0.04 97.37
+w/o summary 15.11 3.56 13.81 0.12 98.34
+w/o qa 12.45 2.83 10.63 0.03 97.27
+NarraSum
+Full T3 25.28 5.88 23.00 2.15 95.91
+w/o summary 26.40 6.34 23.67 2.27 94.86
+w/o qa 23.28 5.14 20.73 1.98 93.69
+GPT-4-turbo NLQuAD
+Full T3 23.66 7.36 21.62 0.36 49.16
+w/o summary 27.78 9.83 27.31 0.64 46.00
+w/o qa 29.78 9.03 25.61 0.12 47.22
+BBC Summary
+Full T3 36.83 14.14 33.27 7.69 69.01
+w/o summary 37.49 16.46 35.97 8.72 63.56
+w/o qa 36.45 15.26 34.07 7.01 62.50
+FairyTaleQA
+Full T3 14.14 3.36 12.84 0.10 97.67
+w/o summary 16.14 3.66 14.44 0.19 97.49
+w/o qa 15.52 3.18 13.14 0.11 96.52
+GPT-4-turbo NarraSum
+Full T3 27.27 7.42 24.45 3.60 97.44
+w/o summary 28.12 5.99 24.90 2.36 95.32
+w/o qa 27.90 5.96 24.02 1.98 94.28
+GPT-4o NLQuAD
+Full T3 31.00 13.03 29.27 1.42 63.95
+w/o summary 36.28 15.29 34.02 3.09 77.85
+w/o qa 41.16 20.29 38.99 5.31 65.23
+BBC Summary
+Full T3 41.71 21.45 39.44 10.94 70.12
+w/o summary 40.17 19.98 38.35 9.13 68.56
+w/o qa 42.13 20.40 39.02 11.06 66.37
+FairyTaleQA
+Full T3 14.86 3.72 13.78 0.07 98.92
+w/o summary 23.80 9.47 24.14 0.23 97.35
+w/o qa 20.60 6.32 19.32 0.08 96.83
+NarraSum
+Full T3 28.86 8.58 25.66 2.55 99.69
+w/o summary 27.44 7.28 24.83 2.45 98.50
+w/o qa 29.03 7.22 26.04 2.71 97.42
+Gemini-1.0-pro NLQuAD
+Full T3 28.11 11.05 26.48 1.25 71.64
+w/o summary 25.50 8.99 23.96 0.42 66.43
+w/o qa 25.21 8.78 23.45 0.45 68.09
+BBC Summary
+Full T3 33.58 13.92 31.17 7.22 55.61
+w/o summary 31.67 11.98 30.06 6.38 45.87
+w/o qa 31.16 10.31 29.85 5.86 42.15
+FairyTaleQA
+Full T3 13.43 2.43 13.39 0.14 94.24
+w/o summary 13.97 3.13 13.07 0.05 96.48
+w/o qa 13.02 2.12 12.18 0.01 95.07
+NarraSum
+Full T3 28.25 5.74 26.36 2.09 91.22
+w/o summary 27.63 5.91 24.59 1.67 93.70
+w/o qa 26.73 5.47 24.06 2.08 92.57
+Gemini-1.5-pro NLQuAD
+Full T3 24.19 7.97 22.44 0.47 65.09
+w/o summary 25.51 8.00 24.48 0.35 68.40
+w/o qa 26.98 8.08 24.93 0.42 67.36
+BBC Summary
+Full T3 35.10 12.58 31.69 6.47 60.49
+w/o summary 37.16 14.76 33.21 7.03 68.21
+w/o qa 37.14 15.77 34.70 9.02 65.94
+FairyTaleQA
+Full T3 17.56 4.34 15.46 0.13 98.53
+w/o summary 14.63 3.24 13.94 0.04 97.74
+w/o qa 14.50 3.20 13.50 0.02 97.20
+Gemini-1.5-pro NarraSum
+Full T3 26.95 5.39 25.32 1.99 93.57
+w/o summary 25.23 5.97 24.37 2.19 93.04
+w/o qa 26.36 6.00 26.04 2.00 92.61
+Claude-3.0-sonnet NLQuAD
+Full T3 33.72 13.27 31.22 2.37 61.69
+w/o summary 35.42 14.30 32.81 2.51 68.94
+w/o qa 33.37 12.20 30.82 3.17 63.14
+BBC Summary
+Full T3 40.90 18.08 37.39 10.74 50.43
+w/o summary 40.94 18.85 36.98 10.06 52.32
+w/o qa 40.27 17.63 36.93 10.09 48.74
+FairyTaleQA
+Full T3 18.91 4.90 17.35 0.33 98.92
+w/o summary 18.24 4.35 16.83 0.24 99.36
+w/o qa 15.21 3.82 14.58 0.08 97.54
+NarraSum
+Full T3 25.87 5.47 22.51 2.05 95.61
+w/o summary 28.34 6.57 26.05 2.47 97.59
+w/o qa 27.52 6.09 25.21 2.02 94.83
+Claude-3.5-sonnet NLQuAD
+Full T3 40.17 18.74 37.92 5.00 82.76
+w/o summary 36.79 16.71 34.45 4.09 75.03
+w/o qa 38.08 15.83 35.80 4.02 72.94
+BBC Summary
+Full T3 43.54 21.13 40.36 13.00 65.83
+w/o summary 40.26 16.83 35.99 13.31 58.32
+w/o qa 41.55 20.72 40.05 31.90 58.57
+FairyTaleQA
+Full T3 20.96 6.07 19.52 0.47 99.97
+w/o summary 19.94 5.84 19.41 0.53 98.53
+w/o qa 19.86 5.36 18.18 0.31 97.98
+NarraSum
+Full T3 29.78 7.79 27.23 3.23 98.14
+w/o summary 29.63 7.60 27.86 2.96 96.43
+w/o qa 29.38 7.43 26.53 2.41 97.61
+A.12 Significance Analysis
+In this analysis, we used the t-test to evaluate the significance of the observed results.
+The null hypothesis $H_{0}$ assumes that there is no significant difference between the two algorithms compared in our experiment.
+The test statistic $t$ is calculated as follows:
+$t=\frac{\bar{x}_{1}-\bar{x}_{2}}{\sqrt{\frac{s_{1}^{2}}{n_{1}}+\frac{s_{2}^{2}}{n_{2}}}}$ (3)
+[…]
+We calculate the sample means $\bar{x}_{1},\bar{x}_{2}$ and compute the sample standard deviations $s_{1},s_{2}$.
+Then we plug values into Equation (3) to determine the test statistic $t$ and compare $t$-value with the critical value from the $t$-distribution at a given confidence level $\alpha=0.05$.
+The p-value for each metric result between our direct summarization and summary with T3 are listed in Table 3. Here we sign the model with * if the p-values for the metric are less than 0.05.
+Models ROUGE-1 ROUGE-2 ROUGE-L BLEU Factscore
+BBC summary
+GPT-4o* 0** 0** 0** 0.0007 0.0004
+GPT-4-turbo 0.0431 0.1714 0.0367 0.0836 0.0227
+GPT-3.5-turbo* 0.0124 0.0388 0.0080 0.0082 0.0017
+Gemini-1.5-pro* 0.0160 0.0089 0.0011 0.0157 0.0160
+Gemini-1.0-pro* 0.0101 0.0191 0.0103 0.0495 0.0091
+Claude-3.5-sonnet* 0.0042 0.0003 0.0043 0.0001 0.0003
+Claude-3.0-sonnet 0.0795 0.1108 0.0292 0.0268 0.5309
+NarraSum
+GPT-4o* 0** 0.0009 0** 0.0407 0.0003
+GPT-4-turbo* 0.0077 0.0114 0.0177 0.0418 0.3652
+GPT-3.5-turbo* 0.0499 0.0431 0.0478 0.0456 0.0203
+Gemini-1.5-pro* 0.0332 0.0074 0.0196 0.0482 0.0459
+Gemini-1.0-pro 0.0198 0.0312 0.0422 0.0696 0.0086
+Claude-3.5-sonnet 0.0182 0.0417 0.0152 0.1526 0.1040
+Claude-3.0-sonnet* 0.0483 0.0099 0.0523 0.0154 0.0691
+NLQuAD
+GPT-4o* 0.0086 0.0084 0.0072 0.0087 0.0420
+GPT-4-turbo* 0.0189 0.0417 0.0162 0.0488 0.0340
+GPT-3.5-turbo* 0.0299 0.0441 0.0376 0.0276 0.0457
+Gemini-1.5-pro* 0.0436 0.0468 0.0282 0.0261 0.0014
+Gemini-1.0-pro* 0.0479 0.0420 0.0289 0.0181 0.0498
+Claude-3.5-sonnet* 0.0061 0.0003 0.0034 0.0002 0.0212
+Claude-3.0-sonnet* 0.0095 0.0403 0.0427 0.0193 0.0089
+FairytaleQA
+GPT-4o* 0.0083 0.0180 0.0100 0.0194 0.0124
+GPT-4-turbo* 0.0073 0.0375 0.0185 0.0367 0.0356
+GPT-3.5-turbo 0.0409 0.0524 0.0895 0.0451 0.1184
+Gemini-1.5-pro* 0.0473 0.0296 0.0188 0.0298 0.0365
+Gemini-1.0-pro* 0.0168 0.0151 0.0169 0.0355 0.0496
+Claude-3.5-sonnet* 0.0204 0.0031 0.0239 0.0423 0.0225
+Claude-3.0-sonnet* 0.0340 0.0176 0.0338 0.0147 0.0092
+Table 3: Experimental Results of the p-value on BBC summary, NarraSum, NLQuAD and FairytaleQA datasets between the result of each LLM without and with T3. 0** means the p-value is less than $10^{-7}$